# Maximum gap in complex Ginibre matrices

### v1

@@ 5. Proofs of Main Results
(27) $\begin{split}&{d_{\mathrm{KR}}}(\Xi_{n}\cap B,\zeta\cap B)\\ &\leq 2\Big{\{}\int_{B}\mathbb{E}[\Xi_{n}(T_{n,z})]\mathbf{L}(\mathrm{d}z)+\int_{B}\mathbb{E}[\Xi_{n}^{z!}(T_{n,z})]\,\mathbf{L}(\mathrm{d}z)+\mathbf{L}(B)\sup_{z\in B}\mathbb{E}[(\Xi_{n}\Delta\Xi_{n}^{z!})(S_{n,z})]\Big{\}}.\end{split}$
@@
We fix $c>0$, $n\in\mathbb{N}$, and a Borel set $B\subset B_{1}(0)$ with $\sup_{z\in B}|z|<1$. We recall that we drop $c$ from the notation, since our arguments do not rely on the exact value of $c$. We have already proved (5) in Lemma 8, so we turn to (4).
We use (27) and bound the three terms in the curly brackets separately. The integral (with respect to $\mathbf{L}$) of the first term is given by
$\int_{B}\mathbb{E}[\Xi_{n}(T_{n,z})]\mathbf{L}(\mathrm{d}z)$ $=\int_{\sqrt{n}B}\int_{\sqrt{n}T_{n,z}}\mathbb{P}(\xi_{n}^{z!}(B_{r_{n}}(z))=0)\mathbb{P}(\xi_{n}^{w!}(B_{r_{n}}(w))=0)\rho_{n}(z)\rho_{n}(w)\,\mathrm{d}w\,\mathrm{d}z.$
Since $n\mathbb{P}(\xi_{n}^{z!}(B_{r_{n}}(z))=0)\rho_{n}(z)=c$ for all $z\in B$ by the definition of $r_{n}$, this integral equals
$c^{2}|B||T_{n,z}|=\frac{\pi c^{2}|B|(\log n)^{2}}{n}.$
Next, we consider the second term in the curly brackets in (27). By (7),
$\int_{B}\mathbb{E}[\Xi_{n}^{z!}(T_{n,z})]\mathbf{L}(\mathrm{d}z)$
$=\int_{B}\mathbb{E}[\Xi_{n}[\xi_{n}^{z,\varnothing}](T_{n,z}\setminus\{z\})]\mathbf{L}(\mathrm{d}z)$
@@
where $\xi_{n}^{z!,w!}:=(\xi_{n}^{z!})^{w!}$. Since the reduced Palm process of a determinantal process is again determinantal (see (8)), $\xi_{n}^{z!,w!}$ is a determinantal process and therefore NA. Assume without loss of generality that $r_{n}(w)\geq r_{n}(z)$. Then by (10) (applied to indicator functions for the disjoint sets $B_{r_{n}}(z)$ and $B_{r_{n}}(w)\setminus B_{r_{n}}(z)$),
@@
Note that by [40, Theorem 1] there is a coupling $(\xi^{\prime},\xi^{\prime\prime})\sim(\xi_{n}^{z!},\xi_{n}^{z!,w!})$ such that $\xi^{\prime\prime}\subset\xi^{\prime}$ and $|\xi^{\prime}\setminus\xi^{\prime\prime}|\leq 1$. Using this coupling, we bound the right-hand side of (30) by
@@
With the same idea applied to $(\xi_{n},\xi_{n}^{z!})$, we have
$\mathbb{P}(\xi_{n}^{z!}(B_{r_{n}}(z))\leq 1)\mathbb{P}(\xi_{n}^{z!}(B_{r_{n}}(w)\setminus B_{r_{n}}(z))\leq 1)\leq\mathbb{P}(\xi_{n}(B_{r_{n}}(z))\leq 2)\mathbb{P}(\xi_{n}(B_{r_{n}}(w)\setminus B_{r_{n}}(z))\leq 2).$
Fix $\epsilon>0$. We obtain from Lemma 7 and Lemma 8 that
@@
where the second inequality follows from the inclusion $B_{r_{n}(w)/2}(w+\frac{w-z}{2})\subset B_{r_{n}}(w)\setminus B_{r_{n}}(z)$, since $r_{n}(w)\geq r_{n}(z)$ by assumption. Since $\rho_{n}^{(2)}\leq 1/\pi^{2}$, we conclude from (28) and (29) that
$\int_{B}\mathbb{E}[\Xi_{n}^{z!}(T_{n,z})]\mathbf{L}(\mathrm{d}z)\leq\frac{n^{\epsilon-1/16}|\sqrt{n}B||\sqrt{n}T_{n,0}|}{\pi^{2}}=\frac{n^{\epsilon-1/16}(\log n)^{2}|B|}{\pi}.$
Finally, we consider the integrated expected symmetric difference term on the right-hand side of (27). Recall that $S_{n,z}:=B\setminus T_{n,z}$ for $z\in B_{1}(0)$. From the definitions of $\Xi_{n}$ and $\Xi_{n}^{z}$ preceding the proof of Theorem 1, we obtain that almost surely,
(31) $(\Xi_{n}\Delta\Xi_{n}^{z})(S_{n,z})$ $\leq(\xi_{n}\Delta\hat{\xi}_{n}^{\sqrt{n}z,\varnothing})(\sqrt{n}S_{n,z})$
(32) $\quad+\sum_{w\in\xi\cap\hat{\xi}_{n}^{\sqrt{n}z,\varnothing}\cap(\sqrt{n}S_{n,z})}\big{|}\mathds{1}\{\hat{\xi}_{n}^{\sqrt{n}z,\varnothing}(B_{r_{n}}(w))=0\}-\mathds{1}\{\xi_{n}(B_{r_{n}}(w))=0\}\big{|}.$
We begin by bounding the expectation of the right-hand side of (31). Since $\xi_{n}^{\sqrt{n}z!}\subset\xi_{n}$, and $\xi_{n}^{\sqrt{n}z!}\subset\hat{\xi}_{n}^{\sqrt{n}z!,\varnothing}$ a.s. on $B_{r_{n}}(\sqrt{n}z)^{c}$, and $B_{r_{n}}(\sqrt{n}z)\subset\sqrt{n}T_{n,z}$, we have
$\mathbb{E}[(\xi_{n}\Delta\hat{\xi}_{n}^{\sqrt{n}z,\varnothing})(\sqrt{n}S_{n,z})]$ $\leq\mathbb{E}[(\xi_{n}\setminus\xi_{n}^{\sqrt{n}z!})(\sqrt{n}S_{n,z})]+\mathbb{E}[(\hat{\xi}_{n}^{\sqrt{n}z!,\varnothing}\setminus\xi_{n}^{\sqrt{n}z!})(\sqrt{n}S_{n,z})]$
@@
(34) $\quad+\mathbb{E}[\hat{\xi}_{n}^{\sqrt{n}z!,\varnothing}(\sqrt{n}S_{n,z})]-\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})].$
Using Lemma 3, (8), and $K_{n}(z,z)=\rho_{n}(z)\geq\frac{1}{2\pi}$ uniformly for $z\in B$ for sufficiently large $n$, the difference in (33) is bounded by
$2\pi\int_{\sqrt{n}S_{n,z}}|K_{n}(\sqrt{n}z,w)|^{2}\,\mathrm{d}w$ $\leq 2\pi|B|n\left(\sup_{w\in\sqrt{n}S_{n,z}}e^{-|w-\sqrt{n}z|^{2}/2}+O(e^{-cn})\right)$
(35) $\leq\frac{2\pi|B|}{{n^{2}}}+O(e^{-cn}),\quad z\in B.$
To bound (34), we use that $\mathbf{L}(\mathrm{d}z)=n\mathbb{P}(\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))=0)\rho_{n}(\sqrt{n}z)=c\,\mathrm{d}z$ for all $z\in B_{1}(0)$ by the definition of $\mathbf{L}$. Together with (6) and (7), this gives for all Borel sets $A\subset B_{1}(0)$ that
@@
$\quad=n\int_{A}\mathbb{E}\Big{[}\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})\mathds{1}\{\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))=0\}\Big{]}\rho_{n}(\sqrt{n}z)\mathrm{d}z.$
Since $A$ is arbitrary, the integrands must agree for $\mathbf{L}$-almost all $z\in B$. Hence, we find that (34) becomes
@@
$\quad={c^{-1}n\rho_{n}(\sqrt{n}z)\mathbb{P}(\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))=0)}\Big{(}\mathbb{E}[\xi_{n}^{\sqrt{n}z!,\varnothing}(\sqrt{n}S_{n,z})]-\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})]\Big{)}$
$\quad={c^{-1}n\rho_{n}(\sqrt{n}z)}\Big{(}\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})\mathds{1}\{\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))=0\}]$
$\hskip 85.35826pt-\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})]\mathbb{P}(\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(z))=0)\Big{)}$
(36) $\quad={c^{-1}n\rho_{n}(\sqrt{n}z)}\mathsf{Cov}(\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z}),\mathds{1}\{\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))=0\}).$
@@
It is easy to see that $f^{(k)},\,k\in\mathbb{N},$ and $f$ are bounded and increasing. Now we use that the reduced Palm process $\xi_{n}^{\sqrt{n}z!}$ is a determinantal process itself and therefore NA (see (10)). Since $r_{n}\leq\log n$ for large enough $n$, by Lemma 8, we obtain
@@
$-c^{-1}n\rho_{n}(\sqrt{n}z)\mathsf{Cov}(\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z}),\xi^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z)))$
(37) $\quad\leq c^{-1}n\rho_{n}(\sqrt{n}z)\big{|}\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})]-\mathbb{E}[\xi_{n}(\sqrt{n}S_{n,z})]\big{|}\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]$
(38) $\qquad+c^{-1}n\rho_{n}(\sqrt{n}z)\big{|}\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]-\mathbb{E}[\xi_{n}(\sqrt{n}S_{n,z})]\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]\big{|}.$
@@
$\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]\leq\mathbb{E}[\xi_{n}(B_{r_{n}}(\sqrt{n}z))]\leq r_{n}(z)^{2}.$
Hence, by Lemma 8 and the bound $\rho_{n}(\sqrt{n}z)\leq\frac{1}{\pi}$,
$c^{-1}n\rho_{n}(\sqrt{n}z)\big{|}\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})]-\mathbb{E}[\xi_{n}(\sqrt{n}S_{n,z})]\big{|}\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]$
(39) $\quad\leq{\frac{4|B|\sqrt{\log n}}{cn}}+O(e^{-cn}),\quad z\in B.$
Next we consider (38). We write $\rho_{n,z}^{(m)}$ for the $m$-th correlation function of $\xi_{n}^{z!}$. By [47, Lemma 6.4],

### v2

@@ 5. Proofs of Main Results
(30) $\begin{split}&{d_{\mathrm{KR}}}(\Xi_{n}\cap B,\zeta_{\kappa}\cap B)\\ &\leq 2\Big\{\int_{B}\mathbb{E}\big[\Xi_{n}(T_{n,z})\big]L_{n}(\mathrm{d}z)+\int_{B}\mathbb{E}\big[\Xi_{n}^{z!}(T_{n,z})\big]\,L_{n}(\mathrm{d}z)+L_{n}(B)\sup_{z\in B}\mathbb{E}\big[(\tilde{\Xi}_{n}\Delta\tilde{\Xi}_{n}^{z!})(S_{n,z})\big]\Big\},\end{split}$
where $T_{n,z}=B_{\log n/\sqrt{n}}(z)$, $S_{n,z}=B\setminus T_{n,z}$, and for any $A\subset\mathbb{C}$,
$\tilde{\Xi}_{n}(A):=\Xi_{n}\left[\xi_{n}|_{B_{r_{n}}(\sqrt{n}z)^{c}}\right](A\cap{S_{n,z}}),\qquad\tilde{\Xi}_{n}^{z!}(A):=\Xi_{n}\left[\xi_{n}^{\sqrt{n}z,\varnothing}|_{B_{r_{n}}(\sqrt{n}z)^{c}}\right](A\cap{S_{n,z}})$
with $\xi_{n}|_{B_{r_{n}}(\sqrt{n}z)^{c}}$ and $\xi_{n}^{\sqrt{n}z,\varnothing}|_{B_{r_{n}}(\sqrt{n}z)^{c}}$ coupled as in Section 3.3.333We recall that the construction of this coupling used Lemma 7, which requires that $n$ be sufficiently large in a way that depends only on $\kappa$ and $g$; this requirement is absorbed into the constant $C(t,\kappa)$ in the statement of this lemma.
Proof.
Let $\mathcal{L}$ be the generator of the Glauber dynamics for the Poisson process $\zeta_{\kappa}$ (see [10, (2.7)]), which is given by
$\mathcal{L}h(\omega)=\int_{\mathbb{C}}D_{z}h(\omega)L_{n}(\mathrm{d}z)-\int_{\mathbb{C}}D_{z}h(\omega-\delta_{z})\omega(\mathrm{d}z),$
where $D_{z}h(\omega)=h(\omega+\delta_{z})-h(\omega)$ and $\mathcal{L}$ acts on functions $h\colon\widehat{\mathbf{N}}\rightarrow\mathbb{R}$. Let $P_{s}$ be the Markov semigroup associated with $\mathcal{L}$. The proof of [10, Theorem 3.1], specialized to our setting, yields
(31) $\begin{split}{d_{\mathrm{KR}}}(\Xi_{n}\cap B,\zeta_{\kappa}\cap B)&\leq\sup_{h\in\text{Lip}}\int_{0}^{\infty}\big|\mathbb{E}[\mathcal{L}P_{s}h(\Xi_{n}\cap B)]\big|\mathrm{d}s\\ &=\sup_{h\in\text{Lip}}\int_{0}^{\infty}\left|\int_{B}\left(\mathbb{E}\big[D_{z}P_{s}h(\Xi_{n}\cap B)\big]-\mathbb{E}\big[D_{z}P_{s}h(\Xi_{n}^{z!}\cap B)\big]\right)L_{n}(\mathrm{d}z)\right|\mathrm{d}s.\end{split}$
In the following we will discuss how to bound the right-hand side of (31). This will involve a coupling outside of a small ball around an arbitrary $z\in B$ and a rough union bound within this ball. By the definition of $D_{z}$, we have
(32) $\begin{split}&\Big|\mathbb{E}\big[D_{z}P_{s}h(\Xi_{n}\cap B)\big]-\mathbb{E}\big[D_{z}P_{s}h(\Xi_{n}^{x!}\cap B)\big]\Big|\\ &=\left|\mathbb{E}\left[P_{s}h(\Xi_{n}\cap B+\delta_{z})-P_{s}h(\Xi_{n}^{z!}\cap B+\delta_{z})-P_{s}h(\Xi_{n}\cap B)+P_{s}h(\Xi_{n}^{z!}\cap B)\right]\right|.\end{split}$
By [10, (2.9)],
$\mathbb{E}\left[\big|P_{s}h(\Xi_{n}\cap B)-P_{s}h(\Xi_{n}\cap S_{n,z})\big|\right]$ $\leq e^{-s}\mathbb{E}\big[\Xi_{n}(T_{n,z})\big],$
$\mathbb{E}\left[\big|P_{s}h(\Xi^{z!}_{n}\cap B)-P_{s}h(\Xi^{z!}_{n}\cap S_{n,z})\big|\right]$ $\leq e^{-s}\mathbb{E}\big[\Xi^{z!}_{n}(T_{n,z})\big],$
$\mathbb{E}\left[\big|P_{s}h(\Xi_{n}\cap B+\delta_{z})-P_{s}h(\Xi_{n}\cap S_{n,z}+\delta_{z})\big|\right]$ $\leq e^{-s}\mathbb{E}\big[\Xi_{n}(T_{n,z})\big],$
$\mathbb{E}\left[\big|P_{s}h(\Xi^{z!}_{n}\cap B+\delta_{z})-P_{s}h(\Xi^{z!}_{n}\cap S_{n,z}+\delta_{z})\big|\right]$ $\leq e^{-s}\mathbb{E}\big[\Xi^{z!}_{n}(T_{n,z})\big].$
Inserting these bounds in (32) gives
(33) $\begin{split}&\Big|\mathbb{E}\big[D_{z}P_{s}h(\Xi_{n}\cap B)\big]-\mathbb{E}\big[D_{z}P_{s}h(\Xi_{n}^{z!}\cap B)\big]\Big|\\ &\leq\left|\mathbb{E}\left[P_{s}h(\Xi_{n}\cap S_{n,z})-P_{s}h(\Xi_{n}^{z!}\cap S_{n,z})\right]\right|\\ &\quad+\left|\mathbb{E}\left[P_{s}h(\Xi_{n}\cap S_{n,z}+\delta_{z})-P_{s}h(\Xi_{n}^{z!}\cap S_{n,z}+\delta_{z})\right]\right|+2e^{-s}\left(\mathbb{E}\big[\Xi_{n}(T_{n,z})\big]+\mathbb{E}\big[\Xi^{z!}_{n}(T_{n,z})\big]\right).\end{split}$
Note that $\Xi_{n}[\omega]\cap S_{n,z}$ depends on $\omega$ only through
$\omega\cap\big(\sqrt{n}S_{n,z}\oplus B_{r_{0}}(0)\big),\qquad r_{0}:=\sup_{w\in\sqrt{n}S_{n,z}}r_{n}(w),$
where $\oplus$ denotes the Minkowski sum. By Lemma 13, there exists $C(t,\kappa)>1$ such that
$\sqrt{n}S_{n,z}\oplus B_{r_{0}}\subset B_{r_{n}(z)}(\sqrt{n}z)^{c}$
for $n\geq C$. Then by [10, (2.9)] and the definitions of $\tilde{\Xi}_{n}$ and $\tilde{\Xi}_{n}^{z!}$ in the statement of the lemma,
(34) $\begin{split}\left|\mathbb{E}\left[P_{s}h(\Xi_{n}\cap S_{n,z})-P_{s}h(\Xi_{n}^{z!}\cap S_{n,z})\right]\right|&\leq e^{-s}\mathbb{E}[(\tilde{\Xi}_{n}\Delta\tilde{\Xi}_{n}^{z!})(S_{n,z})]\leq\sup_{z\in B}e^{-s}\mathbb{E}[(\tilde{\Xi}_{n}\Delta\tilde{\Xi}_{n}^{z!})(S_{n,z})],\\ \left|\mathbb{E}\left[P_{s}h(\Xi_{n}\cap S_{n,z}+\delta_{z})-P_{s}h(\Xi_{n}^{z!}\cap S_{n,z}+\delta_{z})\right]\right|&\leq e^{-s}\mathbb{E}[\tilde{(}\Xi_{n}\Delta\tilde{\Xi}_{n}^{z!})(S_{n,z})]\leq\sup_{z\in B}e^{-s}\mathbb{E}[(\tilde{\Xi}_{n}\Delta\tilde{\Xi}_{n}^{z!})(S_{n,z})].\end{split}$
Inserting this into (33), and integrating over $s$ in (31), we obtain the conclusion. ∎
Remark 15.
In (34), we took suprema over $B$ to ensure that our upper bounds are measurable as functions of the $z\in B$ appearing on the differences on the left-hand sides of the inequalities (and hence can be integrated with respect to $L_{n}(\mathrm{d}z)$). These suprema are constants in $z$, and hence trivially measurable. In contrast, it does not seem apparent from the construction of our couplings that $z\mapsto\mathbb{E}[(\tilde{\Xi}_{n}\Delta\tilde{\Xi}_{n}^{z!})(S_{n,z})]$ is measurable.
We are now ready for the proofs of our main results.
@@
Fix a Borel set $B\subset B_{1}(0)$ with $\sup_{z\in B}|z|<s$. We have already proved (6) in Lemma 13, so we turn to (5).
We use (LABEL:eq:KRbou) and bound the three terms in the curly brackets separately. Using the definitions of $L_{n}$ and $\Xi_{n}$, the first term is
$\int_{B}\mathbb{E}[\Xi_{n}(T_{n,z})]L_{n}(\mathrm{d}z)$ $=\int_{\sqrt{n}B}\int_{\sqrt{n}T_{n,z}}\mathbb{P}(\xi_{n}^{z!}(B_{r_{n}}(z))=0)\mathbb{P}(\xi_{n}^{w!}(B_{r_{n}}(w))=0)\rho_{n}(z)\rho_{n}(w)\,\mathrm{d}w\,\mathrm{d}z.$
Since $n\mathbb{P}(\xi_{n}^{z!}(B_{r_{n}}(z))=0)\rho_{n}(z)=\kappa$ for all $z\in B$ by the definition of $r_{n}$, this integral equals
$\int_{B}\mathbb{E}[\Xi_{n}(T_{n,z})]L_{n}(\mathrm{d}z)=\kappa^{2}|B||T_{n,z}|=\frac{\pi\kappa^{2}|B|(\log n)^{2}}{n}.$
Next, we consider the second term in the curly brackets in (LABEL:eq:KRbou). By Remark 4 and (8),
$\int_{B}\mathbb{E}[\Xi_{n}^{z!}(T_{n,z})]L_{n}(\mathrm{d}z)$
$=\int_{B}\mathbb{E}[\Xi_{n}[\xi_{n}^{z,\varnothing}](T_{n,z}\setminus\{z\})]L_{n}(\mathrm{d}z)$
@@
where we recall from Section 3.3 that $\xi_{n}^{z!,w!}:=(\xi_{n}^{z!})^{w!}$. Since the reduced Palm process of a determinantal process is again determinantal (see (10)), $\xi_{n}^{z!,w!}$ is a determinantal process and therefore has negative associations. Assume without loss of generality that $r_{n}(w)\geq r_{n}(z)$. Then by (15) (applied to indicator functions for the disjoint sets $B_{r_{n}}(z)$ and $B_{r_{n}}(w)\setminus B_{r_{n}}(z)$),
@@
Using the coupling $(\xi_{n}^{z!},\xi_{n}^{z!,w!})$ given in Section 3.3, we bound the right-hand side of (37) by
@@
With the same idea applied to $(\xi_{n},\xi_{n}^{z!})$ (again using the coupling specified in Section 3.3), we have
(38) $\mathbb{P}(\xi_{n}^{z!}(B_{r_{n}}(z))\leq 1)\mathbb{P}(\xi_{n}^{z!}(B_{r_{n}}(w)\setminus B_{r_{n}}(z))\leq 1)\leq\mathbb{P}(\xi_{n}(B_{r_{n}}(z))\leq 2)\mathbb{P}(\xi_{n}(B_{r_{n}}(w)\setminus B_{r_{n}}(z))\leq 2).$
Fix $\epsilon>0$. We obtain from Lemma 12 and Lemma 13 that there exists $C(s,\epsilon,\kappa)>1$ such that for all $n\geq C$,
@@
where the second inequality follows from the inclusion
$B_{r_{n}(w)/2}\left(w+\frac{w-z}{2}\right)\subset B_{r_{n}}(w)\setminus B_{r_{n}}(z),$
since $r_{n}(w)\geq r_{n}(z)$ by assumption. Since $\rho_{n}^{(2)}\leq 1/\pi^{2}$ (by a direct calculation using the explicit form of the kernel), we conclude from (35), (36), (37), and (38) that
$\int_{B}\mathbb{E}[\Xi_{n}^{z!}(T_{n,z})]L_{n}(\mathrm{d}z)\leq\frac{n^{\epsilon-1/16}|\sqrt{n}B||\sqrt{n}T_{n,0}|}{\pi^{2}}=\frac{n^{\epsilon-1/16}(\log n)^{2}|B|}{\pi}.$
Finally, we consider the integrated expected symmetric difference term on the right-hand side of (LABEL:eq:KRbou). Recall that $S_{n,z}:=B\setminus T_{n,z}$ for $z\in B_{1}(0)$. From the definitions of $\tilde{\Xi}_{n}$ and $\tilde{\Xi}_{n}^{z}$ in the statement of Lemma 14 (and the facts noted following (33)), we obtain that almost surely,
(39) $(\tilde{\Xi}_{n}\Delta\tilde{\Xi}_{n}^{z})(S_{n,z})$ $\leq(\xi_{n}\Delta\xi_{n}^{\sqrt{n}z,\varnothing})(\sqrt{n}S_{n,z})$
(40) $\quad+\sum_{w\in\xi_{n}\cap\xi_{n}^{\sqrt{n}z,\varnothing}\cap(\sqrt{n}S_{n,z})}\big|\mathds{1}\{\xi_{n}^{\sqrt{n}z,\varnothing}(B_{r_{n}}(w))=0\}-\mathds{1}\{\xi_{n}(B_{r_{n}}(w))=0\}\big|.$
We begin by bounding the expectation of the right-hand side of (39). From the coupling indicated in Section 3.3, we have
$\xi_{n}^{\sqrt{n}z!}|_{B_{r_{n}}(\sqrt{n}z)^{c}}\subset\xi_{n}|_{B_{r_{n}}(\sqrt{n}z)^{c}},\qquad\xi_{n}^{\sqrt{n}z!}|_{B_{r_{n}}(\sqrt{n}z)^{c}}\subset\xi_{n}^{\sqrt{n}z!,\varnothing}|_{B_{r_{n}}(\sqrt{n}z)^{c}}.$
Using these inclusions in combination with $B_{r_{n}}(\sqrt{n}z)\subset\sqrt{n}T_{n,z}$, we have
$\mathbb{E}[(\xi_{n}\Delta\xi_{n}^{\sqrt{n}z,\varnothing})(\sqrt{n}S_{n,z})]$ $\leq\mathbb{E}[(\xi_{n}\setminus\xi_{n}^{\sqrt{n}z!})(\sqrt{n}S_{n,z})]+\mathbb{E}[(\xi_{n}^{\sqrt{n}z!,\varnothing}\setminus\xi_{n}^{\sqrt{n}z!})(\sqrt{n}S_{n,z})]$
@@
(42) $\quad+\mathbb{E}[\xi_{n}^{\sqrt{n}z!,\varnothing}(\sqrt{n}S_{n,z})]-\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})].$
Using Lemma 8, (10), and $K_{n}(z,z)=\rho_{n}(z)\geq\frac{1}{2\pi}$ uniformly for all $|z|<s$ for $n\geq C(s)$, we can bound the difference in (41) for every $z\in B$ by
$\mathbb{E}[\xi_{n}(\sqrt{n}S_{n,z})]-\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})]$ $\leq 2\pi\int_{\sqrt{n}S_{n,z}}|K_{n}(\sqrt{n}z,w)|^{2}\,\mathrm{d}w$
$\leq 2\pi|B|n\left(\sup_{w\in\sqrt{n}S_{n,z}}e^{-|w-\sqrt{n}z|^{2}/2}+O(e^{-cn})\right)$
(43) $\leq\frac{2\pi|B|}{{n^{2}}}+O(e^{-cn}).$
To bound (42), we use that $L_{n}(\mathrm{d}z)=\kappa\,\mathrm{d}z$ for all $z\in B_{0}(s)$ and $n\geq C(s,\kappa)$ by the definition of $L_{n}$ and Lemma 5. Together with (7) and (8), this gives for all Borel sets $A\subset B_{1}(0)$ that
@@
$\quad=n\int_{A}\mathbb{E}\Big[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})\mathds{1}\{\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))=0\}\Big]\rho_{n}(\sqrt{n}z)\mathrm{d}z,$
where the final equality follows from a change of variables. Since $A$ is arbitrary, the integrands must agree for $L_{n}$-almost all $z\in B$. Hence, we find that (42) becomes
@@
$\quad={\kappa^{-1}n\rho_{n}(\sqrt{n}z)\mathbb{P}(\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))=0)}\Big(\mathbb{E}[\xi_{n}^{\sqrt{n}z!,\varnothing}(\sqrt{n}S_{n,z})]-\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})]\Big)$
$\quad={\kappa^{-1}n\rho_{n}(\sqrt{n}z)}\Big(\mathbb{E}\big[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})\mathds{1}\{\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))=0\}\big]$
$\qquad\qquad\qquad\qquad\qquad-\mathbb{E}\big[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})\big]\mathbb{P}(\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(z))=0)\Big)$
(44) $\quad={\kappa^{-1}n\rho_{n}(\sqrt{n}z)}\mathsf{Cov}\Big(\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z}),\mathds{1}\big\{\xi_{n}^{\sqrt{n}z!}\big(B_{r_{n}}(\sqrt{n}z)\big)=0\big\}\Big).$
@@
It is easy to see that $f^{(k)},\,k\in\mathbb{N},$ and $f$ are bounded and increasing. Now we use that the reduced Palm process $\xi_{n}^{\sqrt{n}z!}$ is itself a determinantal process and therefore has negative associations (see (15)). Since $r_{n}\leq\log n$ for $n\geq C(s,\kappa)$, by Lemma 13, we obtain for such $n$ that
@@
$-\kappa^{-1}n\rho_{n}(\sqrt{n}z)\mathsf{Cov}(\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z}),\xi^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z)))$
(45) $\quad\leq\kappa^{-1}n\rho_{n}(\sqrt{n}z)\big|\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})]-\mathbb{E}[\xi_{n}(\sqrt{n}S_{n,z})]\big|\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]$
(46) $\qquad+\kappa^{-1}n\rho_{n}(\sqrt{n}z)\big|\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]-\mathbb{E}[\xi_{n}(\sqrt{n}S_{n,z})]\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]\big|.$
@@
$\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]\leq\mathbb{E}[\xi_{n}(B_{r_{n}}(\sqrt{n}z))]\leq r_{n}(z)^{2},$
where the first inequality follows from (14) and the second follows from the elementary bound $\rho_{n}\leq\pi^{-1}$. Hence, by Lemma 13 and the bound $\rho_{n}(\sqrt{n}z)\leq\frac{1}{\pi}$, we have using (43) that for all $z\in B$,
$\kappa^{-1}n\rho_{n}(\sqrt{n}z)\big|\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(\sqrt{n}S_{n,z})]-\mathbb{E}[\xi_{n}(\sqrt{n}S_{n,z})]\big|\mathbb{E}[\xi_{n}^{\sqrt{n}z!}(B_{r_{n}}(\sqrt{n}z))]$
(47) $\quad\leq{\frac{4|B|\sqrt{\log n}}{cn}}+O(e^{-cn}).$
Next we consider (46). We write $\rho_{n,z}^{(m)}$ for the $m$-th correlation function of $\xi_{n}^{z!}$. By [50, Lemma 6.4],